After leaving Moose Jaw, Hanson drove
200 kilometers at x km/hr and then decreased his speed by
20 km/hr and drove 240 additional kilometers. Write a rational expression for
his total traveling time. Evaluate the expression for x = 100.

Answer by stanbon(75887)

After leaving Moose Jaw, Hanson drove
200 kilometers at x km/hr and then decreased his speed by
20 km/hr and drove 240 additional kilometers. Write a rational expression for
his total traveling time. Evaluate the expression for x = 100.
-------------------
time = distance/rate
----------------------
Equation:
Total time = 200/x + 240/20
Total time = 200/x + 12
T(100) = 200/100 + 12
T(100) = 2 + 12 = 14 hours.
=================================
Cheers,
Stan H.